Crockett Corporation's 5-year bonds yield 6.35%, and 5-year T-bonds yield 4.45%. The real risk-free rate is r* = 2.80%, the default risk premium for Crockett's bonds is DRP = 1.00% versus zero for T-bonds, the liquidity premium on Crockett's bonds is LP = 0.90% versus zero for T-bonds, and the maturity risk premium for all bonds is found with the formula MRP = (t – 1) × 0.1%, where t = number of years to maturity. What inflation premium (IP) is built into 5-year bond yields?
Answer #1

For a Treasury Security, Nominal Yield Real risk free rate Inflation Premium Maturity Risk Premium where Inflation premium isMaturity risk premium = (t - 1) * 0.1% = (5 - 1) * 0.1% = 0.4%

4.45% = 2.80% + Inflation premium + 0.4%

Inflation premium = 1.25%